OpenStudy (anonymous):

Tom receives allowance every week. He spends three-fourths of his allowance on food. Last week, Tom spent nine dollars on food. What is Tom’s weekly allowance? Write the verbal model, define the variable, write the equation and find the solution. Show your work.

OpenStudy (anonymous):

3/4x=9 3/4 Because he spends 3/4 of his allowance so x is his allowance they tell you he spent 9 dollars. 4/3(9)=x 12=x so his allowance is 12 dollars you can check 12*(3/4)=9

OpenStudy (anonymous):

hope its understandable

OpenStudy (anonymous):

thx ill give medal and I am going to post one more question not here but a new post

OpenStudy (anonymous):

okay i may be off in a couple minutes my class period is almost over good luck

OpenStudy (anonymous):

k thx